# Magnetic Polarisability of Octet Baryons via Lattice QCD

### v2

@@ X Proton-Neutron Magnetic Polarisability Difference
The difference between the magnetic polarisability of the proton and neutron can provide a test of Reggeon dominance Gasser et al. (2015, 2020). Under the assumption of Reggeon dominance, chiral perturbation theory and Baldin sum rules can predict the difference of magnetic polarisabilities.
@@
Figure 15: The difference between the proton and neutron magnetic polarisabilities. This work’s finite-volume and electro-quenched corrected lattice QCD results (FV & EQ Corr.) are extrapolated to the infinite volume (Inf. Vol.) physical point through chiral extrapolation. These results are compared to the experimental results of the PDG Workman and Others (2022), a Reggeon dominance prediction Gasser et al. (2020) and a previous lattice QCD calculation Bignell et al. (2020) which are horizontally offset at the physical pion mass for clarity.
@@
We calculate the difference between the magnetic polarisability of the proton and neutron by construction of a correlation function ratio analogous to Eq. VI, which provides direct access to the polarisability difference
@@
$\delta E_{\beta_{p}-\beta_{n}}$ is fit using the techniques discussed in Sec. VI. We then extrapolate to the physical regime using the formalism discussed in Sec. VIII. When taking the polarisability difference, the $u-d$ symmetry in the leading loop-integral coefficients of the chiral expansion in full QCD causes the chiral contributions to cancel. As such, the extrapolation becomes a simple linear extrapolation. The resulting value at the physical point is
@@
a much better estimate for the difference between the two polarisabilities. Here the statistical uncertainty is given in parentheses and the systematic uncertainty is negligible. Fig. 15 shows the extrapolation to the physical regime and includes the PDG value Workman and Others (2022), a result derived using Reggeon dominance Gasser et al. (2020) and a previous lattice QCD calculation Bignell et al. (2020).
@@ XI Conclusion
A generalised expression for the magnetic polarisability was derived in a simple constituent quark model. Once corrected, this simple model provides accurate predictions for the octet-baryon magnetic polarisabilities examined herein. It also provides the ability to identify key characteristics of the magnetic polarisability and deep insight into the physics that drives the observed patterns. In particular, we identified the importance of opposite-charge quark flavours in generating large magnetic polarisability values. Here the importance of the up quark is manifest. We also focused on the octet-decuplet mass splitting and associated hyperfine interactions that govern the magnitude of the magnetic polarisability transition term.
@@
Turning our attention to lattice simulation techniques, we investigated the behaviour of the underlying correlation functions associated with the correlation function ratio required to extract the magnetic polarisability in the background field method. We saw that in many cases, the correlation function ratio exhibits plateau-like behaviour in spite of the underlying correlation functions not yet having reached single-state isolation. We developed new methods to address this, ensuring the suppression of excited state contamination, and ensuring the uncertainties in our fitted mass shifts were not underestimated.
@@
We also implemented a weighted averaging method to systematically extract the magnetic polarisability from the associated energy shifts obtained from candidate fit windows. This is especially important in cases where different fit windows introduce variation in the fitted values of the polarisability energy shifts. These improved analysis techniques allow the extraction of magnetic polarisability values for the outer octet baryons at a variety of pion masses with unprecedented precision.
@@
The systematics of the lattice QCD calculation were then addressed. Drawing on chiral effective field theory and accounting for both electro-quenching and finite volume effects we extrapolated our simulation results to the physical point. This process produces results that compare favourably with the experimental values for the proton and neutron. Excellent agreement with the precise experimental value for the proton is observed and our prediction for the neutron is much more precise.
@@
Comparison of phenomenological values for the hyperons indicate our results are very precise on the scale of variation in QCD models. Finally a new precise calculation of the difference in the proton and neutron magnetic polarisabilities has been presented.
@@
We have revealed complex dynamics underpinning the magnetic polarisabilities of octet baryons. It would be interesting to examine these dynamics at a more microscopic level, where the quark mass dependence and environment dependence of individual quark sector contributions to the baryon polarisabilities are exposed. The techniques of the current presentation are flexible enough to admit such an analysis and we anticipate reporting on this in the near future.
Another challenge facing the community is gaining access to the light quark-mass regime. Using the techniques presented here, we are unable to resolve a signal of statistical interest. There are several issues that can lead to large uncontrolled statistical fluctuations and it will be important to examine this further to determine the cause of the fluctuations and perhaps develop new techniques to enable an exploration of the chiral regime.
@@
Finally, the $uds$ members of the baryon octet, the $\Lambda$ and $\Sigma^{0}$ remain to be examined. As highlighted in the introduction, a perturbative calculation Wilcox and Lee (2021); Lee et al. (2023) can offer some important advantages in obtaining a clearer understanding of the magnetic polarisabilities of these baryons.
@@ Acknowledgements.
We thank the PACS-CS Collaboration for making their $2+1$ flavour configurations available and the ongoing support of the International Lattice Data Grid (ILDG). Baryon correlation functions were constructed using the COLA software library, developed at the University of Adelaide Kamleh (2023). WK was supported by the Pawsey Supercomputing Centre through the Pawsey Centre for Extreme Scale Readiness (PaCER) program. RB is grateful for support via STFC grant ST/T000813/1 and acknowledges support from a Science Foundation Ireland Frontiers for the Future Project award with grant number SFI-21/FFP-P/10186 (data not available). This work was supported with supercomputing resources provided by the Phoenix HPC service at the University of Adelaide. This research was undertaken with the assistance of resources from the National Computational Infrastructure (NCI), which is supported by the Australian Government. This research is supported by Australian Research Council through Grants No. DP190102215 and No. DP210103706.
@@ Appendix A Chiral coefficients
As discussed in Sec. VIII, calculating a correction for electro-quenching requires identification of the various valence-valence, valence-sea and sea-sea contributions to the magnetic polarisability. These contributions may be written in terms of the SU(3) chiral coefficients. The coefficients $\chi^{2}_{MB}$ are derived by considering an appropriate chiral Lagrangian. The reader is directed to Ref. Shanahan et al. (2013) for a convenient catalogue of these coefficients. The coefficients are written in terms of the SU(6) values $D,F,\mathcal{C}$. We take $D+F=g_{A}=1.267$, $F=\frac{2}{3}D$, $\mathcal{C}=-1.52$ Bignell et al. (2020).
The chiral coefficients for standard SU(3) axial transitions Shanahan et al. (2013) are listed in Tab. 8 and Tab. 9. Note, our definition of the coefficients in the loop integrals differ from Ref. Shanahan et al. (2013) such that their octet coefficients have been squared and divided by 2 and their decuplet coefficients have been squared and multiplied by $4/3$.

### v3

@@ X Proton-Neutron Magnetic Polarizability Difference
The difference between the magnetic polarizability of the proton and neutron can provide a test of Reggeon dominance Gasser et al. (2015, 2020). Under the assumption of Reggeon dominance, chiral perturbation theory and Baldin sum rules can predict the difference of magnetic polarizabilities.
@@
Figure 15: The difference between the proton and neutron magnetic polarizabilities. This work’s finite-volume- and electroquenched-corrected lattice QCD results are extrapolated to the infinite volume physical point through chiral extrapolation. These results are compared to the experimental results of the PDG Workman and Others (2022), a Reggeon dominance prediction Gasser et al. (2020), and a previous lattice QCD calculation Bignell et al. (2020) which are horizontally offset at the physical pion mass for clarity.
@@
We calculate the difference between the magnetic polarizability of the proton and neutron by construction of a correlation-function ratio analogous to Eq. VI, which provides direct access to the polarizability difference
@@
+($\phantom{=}+\order*{B^{3}}.$ ) . ((95))
$\delta E_{\beta_{p}-\beta_{n}}$ is fit using the techniques discussed in Sec. VI. We then extrapolate to the physical regime using the formalism discussed in Sec. VIII. When taking the polarizability difference, the $u-d$ symmetry in the leading loop-integral coefficients of the chiral expansion in full QCD causes the chiral contributions to cancel. As such, the extrapolation becomes a simple linear extrapolation. The resulting value at the physical point is
@@
a much better estimate for the difference between the two polarizabilities. Here the statistical uncertainty is given in parentheses and the systematic uncertainty is negligible. Figure 15 shows the extrapolation to the physical regime and includes the PDG value Workman and Others (2022), a result derived using Reggeon dominance Gasser et al. (2020) and a previous lattice QCD calculation Bignell et al. (2020).
@@ XI Conclusion
A generalized expression for the magnetic polarizability was derived in a simple constituent quark model. Once scaled, this simple model provides accurate predictions for the octet-baryon magnetic polarizabilities examined herein. It also provides the ability to identify key characteristics of the magnetic polarizability and deep insight into the physics that drives the observed patterns. In particular, we identified the importance of opposite-charge quark flavors in generating large magnetic polarizability values. Here the importance of the up quark is manifest. We also focused on the octet-decuplet mass splitting and associated hyperfine interactions that govern the magnitude of the magnetic polarizability transition term.
@@
Turning our attention to lattice simulation techniques, we investigated the behavior of the underlying correlation functions associated with the correlation-function ratio required to extract the magnetic polarizability in the background-field method. We saw that in many cases, the correlation-function ratio exhibits plateaulike behavior in spite of the underlying correlation functions not yet having reached single-state isolation. We developed new methods to address this, ensuring the suppression of excited-state contamination, and ensuring the uncertainties in our fitted mass shifts were not underestimated.
@@
We also implemented a weighted averaging method to systematically extract the magnetic polarizability from the associated energy shifts obtained from candidate fit windows. This is especially important in cases where different fit windows introduce variation in the fitted values of the polarizability energy shifts. These improved analysis techniques allow the extraction of magnetic polarizability values for the outer octet baryons at a variety of pion masses with unprecedented precision.
@@
The systematics of the lattice QCD calculation were then addressed. Drawing on chiral effective field theory and accounting for both electroquenching and finite volume effects we extrapolated our simulation results to the physical point. This process produces results that compare favorably with the experimental values for the proton and neutron. Excellent agreement with the precise experimental value for the proton is observed and our prediction for the neutron is much more precise.
@@
Comparison of phenomenological values for the hyperons indicate our results are very precise on the scale of variation in QCD models. Finally, a new precise calculation of the difference in the proton and neutron magnetic polarizabilities has been presented.
@@
We have revealed complex dynamics underpinning the magnetic polarizabilities of octet baryons. It would be interesting to examine these dynamics at a more microscopic level, where the quark mass dependence and environment dependence of individual quark sector contributions to the baryon polarizabilities are exposed. The techniques of the current presentation are flexible enough to admit such an analysis and we anticipate reporting on this in the near future.
Another challenge facing the community is gaining access to the light-quark mass regime. Using the techniques presented here, we are unable to resolve a signal of statistical interest for the lightest PACS-CS ensemble. We are currently exploring the possibility that the electroquenching of the sea-quark sector is creating improbable gauge fields when the light-quark ensemble is used at finite magnetic-field strength. Filtering techniques to identify exceptional configurations are under development and the veracity of the approach is under investigation.
@@
Finally, the $uds$ members of the baryon octet, the $\Lambda$, and $\Sigma^{0}$ remain to be examined. As highlighted in the Introduction, a perturbative calculation Wilcox and Lee (2021); Lee et al. (2023) can offer some important advantages in obtaining a clear understanding of the magnetic polarizabilities of these baryons.
The results presented herein are founded on the PACS-CS ensembles Aoki et al. (2009). The ensembles are created by the PACS-CS collaboration, Aoki et al. and are available from jld . The ensembles in question are part of the April 2009 release.
@@ Acknowledgements.
It is a pleasure to acknowledge beneficial communications with Harald Griesshammer. We thank the PACS-CS Collaboration for making their $2+1$ flavor configurations available and the ongoing support of the International Lattice Data Grid (ILDG). Baryon correlation functions were constructed using the COLA software library, developed at the University of Adelaide Kamleh (2023). WK was supported by the Pawsey Supercomputing Centre through the Pawsey Centre for Extreme Scale Readiness (PaCER) program. RB is grateful for support via STFC grant ST/T000813/1 and acknowledges support from a Science Foundation Ireland Frontiers for the Future Project award with grant number SFI-21/FFP-P/10186 (data not available). This work was supported with supercomputing resources provided by the Phoenix HPC service at the University of Adelaide. This research was undertaken with the assistance of resources from the National Computational Infrastructure (NCI), which is supported by the Australian Government. This research is supported by Australian Research Council through Grants No. DP190102215 and No. DP210103706.
@@ Appendix A Chiral coefficients
As discussed in Sec. VIII, calculating a correction for electroquenching requires identification of the various valence-valence, valence-sea and sea-sea contributions to the magnetic polarizability. These contributions may be written in terms of the SU(3) chiral coefficients. The coefficients $\chi^{2}_{MB}$ are derived by considering an appropriate chiral Lagrangian. The reader is directed to Ref. Shanahan et al. (2013) for a convenient catalogue of these coefficients. The coefficients are written in terms of the SU(6) values $D,F,\mathcal{C}$. We take $D+F=g_{A}=1.267$, $F=\frac{2}{3}D$, $\mathcal{C}=-1.52$ Bignell et al. (2020).
The chiral coefficients for standard SU(3) axial transitions Shanahan et al. (2013) are listed in Table 8 and Table 9. Note, our definition of the coefficients in the loop integrals differ from Ref. Shanahan et al. (2013) such that their octet coefficients have been squared and divided by 2 and their decuplet coefficients have been squared and multiplied by $4/3$.